OpenStudy (phebe):

What equation is the equation of a line that passes through (-10,3) ad is perpendicular to y=5x-7

OpenStudy (phebe):

want answer choices

OpenStudy (anonymous):

Yes please

OpenStudy (phebe):

ok

OpenStudy (phebe):

y=5x+53 y=-1/5x-7 y=-1/5x+1 y=1/5x+5

ganeshie8 (ganeshie8):

what do you know about the slopes of perpendicular lines ?

OpenStudy (phebe):

nothing really

OpenStudy (festinger):

As a hint, the product of the gradient of line A and the perpendicular line B is -1. Line A: y=mx+c Line B: y=nx+d If A and B are perpendicular, n*m=-1

OpenStudy (phebe):

thanx but i got the answer now